What would be the equation of the ellipse whose vertices are at (4,5) and (10,5) and which has a semi-minor axis length of 1?
Answer by lwsshak3(11628)

What would be the equation of the ellipse whose vertices are at (4,5) and (10,5) and which has a semi-minor axis length of 1?
**
Given ellipse has a horizontal major axis.
Its standard form: (x-h)^2/a^2+(y-k)^2/b^2=1, a>b, (h,k)=(x,y) coordinates of center
y-coordinate of center=5
x-coordnate of center=(10+4)/2=7 (mid point formula)
center: (7,5)
..
length of major axis=6 (4 to 10)=2a
a=3
a^2=9
..
length of minor axis=2=2b
b=1
b^2=1
..
Equation:
(x-7)^2/9+(y-5)^2/1=1
(x-7)^2/9+(y-5)^2=1